40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 5 15:28:38 2014, Steve, Update, SUS, HeNe laser test preparation beforeRSMqpdMoved.jpgafterSRMqpdMoved.jpg1103Ptest.pngND1onQPD.png
    Reply  Fri Sep 5 16:25:36 2014, Steve, Update, SUS, center ITMY oplev 
    Reply  Tue Sep 9 08:22:28 2014, Steve, Update, SUS, HeNe laser test preparation HeNecompared.png
       Reply  Tue Sep 9 16:59:32 2014, Steve, Update, SUS, HeNe laser test preparation 1103PqpdTranss.png
          Reply  Wed Sep 10 14:58:58 2014, Jenne, Update, SUS, OpLev test: New channels 
             Reply  Thu Sep 11 17:12:42 2014, Steve, Update, SUS, OpLev test: old SP qpd connected 
                Reply  Fri Sep 12 16:30:48 2014, Steve, Update, SUS, HeNe laser test preparation ND2.pngqpd222ND21103P.jpg
                   Reply  Wed Sep 17 13:41:00 2014, Jenne, Update, SUS, New calibrated channels for test QPD NewQPDcalibratedChans.png
                      Reply  Thu Sep 18 09:17:47 2014, Steve, Update, SUS, HeNe laser test  calibrated_16hrs.pngQPD222cal.jpg1103P40s40min.png
                         Reply  Mon Sep 22 10:14:07 2014, Steve, Update, SUS, HeNe laser test  1103Pmon3d.png
                            Reply  Tue Sep 23 08:34:14 2014, Steve, Update, SUS, HeNe laser test  tempDrop.png
                               Reply  Thu Sep 25 10:53:32 2014, Steve, Update, SUS, HeNe laser test  centeredDay6.png
          Reply  Wed Sep 10 15:31:16 2014, Steve, Update, SUS, HeNe laser test preparation SRMqpdisBack0nSRMsus.pngslowVSfast.png
    Reply  Wed Oct 15 17:09:49 2014, manasa, Update, General, Diode laser test preparation 
       Reply  Wed Oct 29 18:07:28 2014, manasa, Update, General, Diode laser test preparation 
Message ID: 10489     Entry time: Wed Sep 10 15:31:16 2014     In reply to: 10479
Author: Steve 
Type: Update 
Category: SUS 
Subject: HeNe laser test preparation 

Quote:

Quote:

 

 SRM as set up in Atm4  26,000 count compared with ETMY oplev servo in operation 7,500 counts for 3 days

Next steps: measure beam size at qpd,

                  place qpd on translation stage for calibration,

                  change 1103P mount to single one

 SRM qpd is installed on translation stage and the shims removed from laser V mounts.

The ETMY oplev servo is on.

SRM oplev servo:  100 microrad/count is an estimate, not calibrated one.

 SRM qpd is back to its normal position. The mount base is still on delrin base. SRM and ITMY need centering.

Tomorrow I will set up the HeNe laser test at the SP table with Ontrack qpd

ETMY oplev servo on. SRM qpd with ND1 ------no component------- 1103P

Attachment 1: SRMqpdisBack0nSRMsus.png  367 kB  | Show | Show all
Attachment 2: slowVSfast.png  43 kB  Uploaded Wed Sep 10 16:46:54 2014  | Show | Show all
ELOG V3.1.3-